# DPFlow: Adaptive Optical Flow Estimation with a Dual-Pyramid Framework

### v1

@@ Abstract
H. Morimitsu, X. Zhu, R. M. Cesar-Jr, X. Ji, and X.-C. Yin, “DPFlow: Adaptive Optical Flow Estimation with a Dual-Pyramid Framework” in CVPR, 2025.
@@ 3.1 Dual-Pyramid Encoder
Figure 3: (a) Feature pyramid. (b) Image pyramid. $\phi$ represents a network block, $I$ denotes images and $X$ features.
@@
where CGU Block is our Cross-Gated Unit block, $I$ represents the original input image, and $I_{s}$ is downsampled to the scale $s$. $\phi$ denotes a convolution block and the superscripts $f,b,i$ denote forward, backward, and image, respectively. We initialize the state $H^{b}$ as zero, and the output feature $X_{s}$ is obtained by concatenating the intermediate forward, backward, and image pyramid features. The same process is applied to both images of the input pair.
@@ 4 Kubric-NK Dataset
Optical flow methods require resolution generalization as flow vectors correlate with input spatial resolution. Nonetheless, current optical flow evaluation datasets do not provide a way to evaluate the robustness of optical flow methods to changes in resolution. Most papers [68, 62, 45] show generalization to high resolutions with only a few qualitative examples, which makes it hard to compare different methods. In this paper, we employ Kubric [12] to generate a multi-resolution synthetic optical flow dataset. We procedurally generate 30 sequences containing 600 unique densely annotated samples at four resolutions: 1K, 2K, 4K, and 8K. In addition to the optical flow, other dense annotations, such as depth, surface normals, and object coordinates, are also available. We define the input resolution $n$K as $(960n,540n)$ and thus refer to our dataset as Kubric-NK. Figure 6 shows some samples from the proposed Kubric-NK dataset.
@@ 5.2 High-Resolution Evaluation
Previous works usually only provide qualitative evaluation for generalization to high-resolution inputs [68, 62]. Here, we provide quantitative metrics of current optical flow models on various existing high-resolution samples and our Kubric-NK dataset. Table 1 provides an overview of the datasets we use in our evaluation. The results are collected using open-source implementations of the models [61, 37]. DPFlow automatically chooses the number of pyramid levels $N$ based on the diagonal input size: $N=\text{round}(\log_{2}(\max(1,\sqrt{W^{2}+H^{2}}/\sqrt{960^{2}+540^{2}})))+3$ and adopts 4 refinement iterations per level.
@@ 5.2.1 Evaluation Checkpoint
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 8 Training and evaluation details
Table 9: Hyperparameters for each training stage. Batch sizes ×$\times 4$ indicate that gradients were accumulated during 4 iterations before backpropagating. Dataset abbreviations denote C: FlyingChairs [11], T: FlyingThings3D [34], H: HD1K [25], S: Sintel [5], K: KITTI 2015 [36], Sp: Spring [35].
@@ 10 Model details
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 11.1 Metrics
where []delimited-[][$\cdot$][ ⋅ ] is the Iverson bracket. Each benchmark selects a specific threshold value and represents its metric by a different name. For example, the KITTI benchmark [36] adopts $\tau=\max{(3,0.05\|G_{i}\|_{2})}$ and calls their metric Fl-All. On the other hand, the Spring benchmark [35] uses $\tau=1$ and names their metric as 1px.
@@
where []delimited-[][$\cdot$][ ⋅ ] is the Iverson bracket, $\delta_{i}=\frac{i}{20}$ and $w_{i}=1-\frac{i-1}{100}$.
@@ 14 Input downsampling
[…]
@@
[…]

### v2

@@ Abstract
H. Morimitsu, X. Zhu, R. M. Cesar-Jr, X. Ji, and X.-C. Yin, “DPFlow: Adaptive Optical Flow Estimation with a Dual-Pyramid Framework” in CVPR, 2025. [Online]. DOI: 10.1109/CVPR52734.2025.01659. Available: https://ieeexplore.ieee.org/document/11094780
@@ 3.1 Dual-Pyramid Encoder
Figure 3: (a) Feature pyramid. (b) Image pyramid. $\phi$ represents a network block, II denotes images and $X$ features.
@@
where CGU Block is our Cross-Gated Unit block, II represents the original input image, and $I_{s}$ is downsampled to the scale $s$. $\phi$ denotes a convolution block and the superscripts $f,b,i$ denote forward, backward, and image, respectively. We initialize the state $H^{b}$ as zero, and the output feature $X_{s}$ is obtained by concatenating the intermediate forward, backward, and image pyramid features. The same process is applied to both images of the input pair.
@@ 4 Kubric-NK Dataset
Optical flow methods require resolution generalization as flow vectors correlate with input spatial resolution. Nonetheless, current optical flow evaluation datasets do not provide a way to evaluate the robustness of optical flow methods to changes in resolution. Most papers [68, 62, 45] show generalization to high resolutions with only a few qualitative examples, which makes it hard to compare different methods. In this paper, we employ Kubric [12] to generate a multi-resolution synthetic optical flow dataset. We procedurally generate 30 sequences containing 600 unique densely annotated samples at four resolutions: 1K, 2K, 4K, and 8K. In addition to the optical flow, other dense annotations, such as depth, surface normals, and object coordinates, are also available. We define the input resolution nnK as $(960n,540n)$ and thus refer to our dataset as Kubric-NK. Figure 6 shows some samples from the proposed Kubric-NK dataset.
@@ 5.2 High-Resolution Evaluation
Previous works usually only provide qualitative evaluation for generalization to high-resolution inputs [68, 62]. Here, we provide quantitative metrics of current optical flow models on various existing high-resolution samples and our Kubric-NK dataset. Table 1 provides an overview of the datasets we use in our evaluation. The results are collected using open-source implementations of the models [61, 37]. DPFlow automatically chooses the number of pyramid levels NN based on the diagonal input size: $N=\text{round}(\log_{2}(\max(1,\sqrt{W^{2}+H^{2}}/\sqrt{960^{2}+540^{2}})))+3$ and adopts 4 refinement iterations per level.
@@ 5.2.1 Evaluation Checkpoint
Table 3: Results on high-resolution datasets. In Spring, ft. indicates results obtained after fine-tuning the model with Spring samples, while 0-sh. is the zero-shot evaluation. $\dagger$} results are not available due to running out of memory. denotes methods that use input tiling [19, 15] and methods with adaptive pyramids. The metrics of each dataset are explained in Sec. 5.1 and in the Sup. Mat.
@@
GMA [23] ICCV’21 3.75 $\dagger$} - 7.07 71.7 $\dagger$} 0.66 1.30 $\dagger$} $\dagger$}
DIP [68] CVPR’22 3.64 9.53 - - 71.3 16.7 0.68 1.57 5.12 $\dagger$}
GMFlow [63] CVPR’22 6.71 $\dagger$} - 10.35 56.7 $\dagger$} 0.88 1.38 $\dagger$} $\dagger$}
@@
SKFlow [54] NeurIPS’22 3.79 $\dagger$} - - 71.9 $\dagger$} 0.62 1.07 $\dagger$} $\dagger$}
@@
CroCo-Flow [61] ICCV’23 3.51 $\dagger$} 4.56 - 76.1 $\dagger$} 0.45 1.06 $\dagger$} $\dagger$}
MS-RAFT+ [20] IJCV’23 3.08 $\dagger$} - 5.72 - $\dagger$} 0.55 1.27 $\dagger$} $\dagger$}
GMFlow+ [64] TPAMI’23 5.72 $\dagger$} - - 69.3 $\dagger$} 0.45 0.97 $\dagger$} $\dagger$}
@@
MemFlow [9] CVPR’24 3.24 $\dagger$} 4.48 5.76 74.6 $\dagger$} 0.55 0.97 $\dagger$} $\dagger$}
@@
CCMR+ [21] WACV’24 3.27 $\dagger$} - - 75.2 $\dagger$} 0.54 1.41 $\dagger$} $\dagger$}
@@ 8 Training and evaluation details
Table 9: Hyperparameters for each training stage. Batch sizes $\times 4$ indicate that gradients were accumulated during 4 iterations before backpropagating. Dataset abbreviations denote C: FlyingChairs [11], T: FlyingThings3D [34], H: HD1K [25], S: Sintel [5], K: KITTI 2015 [36], Sp: Spring [35].
@@ 10 Model details
Table 11: Model details. denotes methods that use input tiling [19, 15]. indicates that the model is running in local correlation mode, which decreases memory consumption but increases inference time. $\dagger$} results are not available due to running out of memory. Times are calculated using an NVIDIA RTX3090 GPU.
@@
RAFT [56] 5.10 5.25 0.51 1.36 4.76 $\dagger$} 0.65 1.71 6.40 $\dagger$}
GMA [23] 5.25 5.88 0.18 1.17 $\dagger$} $\dagger$} 1.41 17.3 $\dagger$} $\dagger$}
DIP [68] 4.21 5.37 0.51 2.11 8.60 $\dagger$} 1.59 5.46 20.9 $\dagger$}
SKFlow [54] 4.84 6.27 0.31 1.70 $\dagger$} $\dagger$} 1.17 17.4 $\dagger$} $\dagger$}
@@
MS-RAFT+ [20] 4.19 16.1 0.74 2.82 $\dagger$} $\dagger$} 2.61 9.34 $\dagger$} $\dagger$}
GMFlow+ [64] 4.49 7.36 0.23 1.69 $\dagger$} $\dagger$} 2.99 20.6 $\dagger$} $\dagger$}
@@
SEA-RAFT (L) [60] 4.30 19.6 0.24 0.70 2.60 $\dagger$} 0.79 2.02 7.21 $\dagger$}
CCMR+ [21] 3.86 11.5 1.34 5.36 $\dagger$} $\dagger$} 3.19 12.1 $\dagger$} $\dagger$}
@@ 11.1 Metrics
where $\cdot$] is the Iverson bracket. Each benchmark selects a specific threshold value and represents its metric by a different name. For example, the KITTI benchmark [36] adopts $\tau=\max{(3,0.05\|G_{i}\|_{2})}$ and calls their metric Fl-All. On the other hand, the Spring benchmark [35] uses $\tau=1$ and names their metric as 1px.
@@
where $\cdot$] is the Iverson bracket, $\delta_{i}=\frac{i}{20}$ and $w_{i}=1-\frac{i-1}{100}$.
@@ 14 Input downsampling
Table 16: Results on high-resolution benchmarks with and without downsampling the inputs. The @1K columns indicate that the inputs are downsampled to 1K resolution and the predictions are upsampled back to the original sizes. $\dagger$} results are not available due to running out of memory.
@@
GMA [23] 3.75 4.53 $\dagger$} 8.23 $\dagger$} 4.91
DIP [68] 3.64 3.87 16.7 8.01 $\dagger$} 5.42
SKFlow [54] 3.79 4.20 $\dagger$} 6.88 $\dagger$} 4.72